A certain gas at 101.325 kPa and 16°C whose volume is 2.83 m³ are compressed into a storage vessel of 0.31 m³ capacity. Before admission, the storage vessel contained the gas at a pressure and temperature of 137.8 kPa and 24°C; after admission the pressure has increased to 1171.8 kPa. What should be the final temperature of the gas in the vessel in Kelvin? 


1
Expert's answer
2021-10-19T15:13:45-0400

"P_1= 101.325\textsf{kPa}\\
T_1 = 16°C = (16+273)\textsf{ K} = 289\textsf{ K}\\
V _1= 2.83\textsf{ m³}\\
P_2= 1171.8\textsf{ kPa}\\
V_2=V_3= 0.31\textsf{ m³}\\
P_3= 137.8 \textsf{ kPa}\\
T_3 = 24°C = (24+273)\textsf{ K} = 297\textsf{ K}"


m of gas to be compressed in the storage vessel

"PV = m_1RT\\m3 = m 2 + m 1"


"(101.325)(2.83)= m R(16+ 273)"

m of gas in storage"(m_1 )= 0.99221\/R"


for final temperature, T3:

"PV = mRT_3\\
(1171.8)(0.31)= PV/T_3= 1.13604 \\T_3=
319.758°K"